OpenStudy (anonymous):

(4x+10)^1/4=(9+7x)^1/4

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

maybe just look it up

hero (hero):

Quad both sides: \[\sqrt[4]{(4x+10)^{1/4}} = \sqrt[4]{(9+7x)^{1/4}}\] Quad and quad root cancels, leaving just: \[4x + 10 = 9 + 7x\]

hero (hero):

@kypaige1